Web Cites Research projects a rate of return of 20% on new projects. Management plans to plow back 30% of all earnings into the firm. Earnings this year will be $3 per share, and investors expect a 12% rate of return on stocks facing the same risks as Web Cites.

a. What is the sustainable growth rate? (Enter your answer as a whole percent.)

b. What is the stock price? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

c. What is the present value of growth opportunities (PVGO)? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

d. What is the P/E ratio? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

e. What would the price and P/E ratio be if the firm paid out all earnings as dividends? (Round your answers to 2 decimal places.)
Answer #1

Answer to Part a.

Sustainable Growth rate (g) = Required Return * Plowback Ratio
Sustainable Growth rate (g) = 0.20 * 0.30
Sustainable Growth rate (g) = 6%

Answer to Part b.

Earnings per Share = $3.00
Dividend Payout Ratio = 1 - Plowback Ratio
Dividend Payout Ratio = 1 – 0.30 = 0.70 or 70%

Dividend per Share = $3.00 * 70% = $2.10

Current Stock Price (P0) = $2.10 * (1.06) / (0.12 – 0.06)
Current Stock Price (P0) = $37.10

Answer to Part c.

No Growth Value = $2.10 / 0.12
No Growth Value = $17.50

Present Value of Growth Opportunities (PVGO) = $37.10 - $17.50
Present Value of Growth Opportunities (PVGO) = $19.60

Answer to Part d.

Price Earnings Ratio = Price per Share / Earnings per Share
Price Earnings Ratio = $37.10 / $3.00
Price Earnings Ratio = 12.37

Answer to Part e.

Current Stock Price = $3.00 / 0.12
Current Stock Price = $25.00

Price Earnings Ratio = Price per Share / Earnings per Share
Price Earnings Ratio = $25.00 / $3.00
Price Earnings Ratio = 8.33